# The Boundaries of Tractability in Hierarchical Task Network Planning

### v1

@@ 4 Primitive Task Networks
Let $\max_{r}(a)$ and $\min_{r}(a)$ be defined as above, but for the reduced instance. Hence, $\max_{r}(a_{1})=\max(a_{1})+\max(a_{2})$ and $\min_{r}(a_{1})=\min(a_{1})+\min(a_{2})$. Thus, any task sequence that solves the original instance also solves the reduced instance. Now assume there is a task sequence that solves the reduced instance. If the same task sequence solves the original instance, we are done. Otherwise, let $x_{1}$ and $x_{2}$ be the number of occurrences of $a_{1}$- and $a_{2}$-tasks, respectively, in the sequence resolving the reduced instance according to the original mapping $\alpha$. Note that $\max_{r}(a_{1})=\max(a_{1})+\max(a_{2})\geq x_{1}+x_{2}\geq\min(a_{1})+\min(a_{2})$. Due to the equivalence in their actions and neighborhoods in $\prec^{+}$, $a_{1}$- and $a_{2}$-tasks can be used interchangeably in the sequence. Thus, any $a_{1}$-task can be replaced by an $a_{2}$-task, and vice versa, which, due to the above inequalities, makes it possible to find a valid solution for the original instance, i.e., one in which $\max(a_{1})\geq x_{1}\geq\min(a_{1})$ and $\max(a_{2})\geq x_{2}\geq\min(a_{2})$. ∎
@@
Note that an instance of Action Executability with $|S|>|T|$ is a trivial no-instance, so we assume $|S|\leq|T|$. By combining the above reduction rules with layered branching techniques and integer programming, we prove:
Theorem 10.
Action Executability is polynomial-time solvable on primitive antichain task networks with constantly many states.
Proof.
@@
We begin by exhaustively applying Reduction Rules R0 and R1, which takes at most $\mathcal{O}(|{\cal I}|^{3})$ time. Afterwards, we have either correctly solved the instance, or—since the network is an antichain—obtained a new equivalent instance where the number of actions is bounded by a function of the number $k$ of states, and in particular $|A|\leq(k+1)^{k}$.
@@
The algorithm now branches over the at most $2^{k}\cdot|A|^{k}$ simple edge paths in ${\cal G}$ that start at $s_{0}$. Let $C$ be the set of distinct simple edge cycles in ${\cal G}$, where two cycles are considered distinct if and only if they do not contain the exact same vertices, i.e., cycles obtained from other cycles by inverting their directions are not counted as different cycles. Then, $|C|\leq 2^{k}\cdot|A|^{k}$. The algorithm branches on all possible subsets $C^{\prime}\subseteq C$ to decide which of these cycles are used at least once by a solution sequence. To ensure that there is a walk, set all cycles in $C^{\prime}$ as unmarked and exhaustively mark cycles that either share a vertex with the chosen path or an already marked cycle. Proceed with $C^{\prime}$ only if all cycles are marked by this procedure.
@@
We write an Integer Linear Program (ILP) as follows. Let the ILP have a variable $x_{c}$ for every $c\in C^{\prime}$. For each action $a$ and $c\in C^{\prime}$, let $c(a)$ denote the number of occurrences of $a$ in $c$. Further, let $p(a)$ denote the number of occurrences of $a$ in the chosen path. For each $a\in A$, require the two constraints
$p(a)+c_{1}(a)x_{c_{1}}+c_{2}(a)x_{c_{2}}+\dots c_{|C^{\prime}|}(a)x_{c_{|C^{\prime}|}}$ $\leq\max(a)$
$p(a)+c_{1}(a)x_{c_{1}}+c_{2}(a)x_{c_{2}}+\dots c_{|C^{\prime}|}(a)x_{c_{|C^{\prime}|}}$ $\geq\min(a),$
@@
where $c_{1},\ldots,c_{|C^{\prime}|}$ are the elements of $C^{\prime}$. Further, for all $c\in C^{\prime}$, add a constraint $x_{c}\geq 1$ to ensure that each of the selected cycles is walked at least once. The algorithm decides that the given instance is a yes-instance if and only if there is a solution to the ILP.
@@
Computing $\min(a)$ and $\max(a)$ for all $a\in A$ takes $\mathcal{O}\!\left(|T|+|S|\right)$ time. Combining the choices for the path and $C^{\prime}$ gives at most $2^{k}\cdot|A|^{k}\cdot 2^{2^{k}\cdot|A|^{k}}$ branches. Testing a set $C^{\prime}$ by the marking procedure takes $\mathcal{O}\!\left(|C|^{2}k\right)$ time, but this time is dominated by the time to solve the ILP. The ILP has at most $|C|\leq 2^{k}\cdot|A|^{k}$ variables and $2|A|+|C|={\mathcal{O}\!\left(2^{k}\cdot|A|^{k}\right)}$ constraints, in which the absolute values of all coefficients are bounded by $|T|$.
By the result of Lenstra (1983) and its subsequent improvements Kannan (1987); Frank and Tardos (1987) an ILP instance ${\cal I}_{\text{ILP}}$ with $n$ variables can be solved in $\mathcal{O}\!\left(n^{O(n)}\cdot|{\cal I}_{\text{ILP}}|\right)$ time. We note that, for the given ILP, $|I_{\text{ILP}}|={\mathcal{O}\!\left(2^{k}\cdot|A|^{k}\cdot\log(|T|)\right)}$. Thus, solving the ILP takes $\mathcal{O}\!\left((2^{k}\cdot|A|^{k})^{O(2^{k}\cdot|A|^{k})}\cdot 2^{2k}\cdot|A|^{2k}\cdot\log(|T|)\right)$ time. Hence, the total runtime can be upper-bounded by
@@
Recalling the bound on $|A|$ argued in the first paragraph of the proof, we obtain that any instance of Action Executability in which the task network is primitive and $\prec^{+}=\emptyset$ can be solved in time $\mathcal{O}(f(k)\cdot\log(|{\cal I}|)+|{\cal I}|^{3})$, for some computable function $f$ of the number $k$ of states. ∎
Using Generalized Partial Order Width.
At this point, we have seen that all of the considered problems on HTNs become polynomial-time solvable on chains and antichains. In this subsection, we unify and extend these positive results by establishing tractability for all HTNs of constant generalized partial order width (whereas chains and antichains have a generalized partial order width of $1$ and 00, respectively). Intuitively speaking, the following three theorems are obtained by “supercharging” the proof techniques used for chains and antichains via a careful dynamic programming routine.
Theorem 11.
Plan Validation is polynomial-time solvable on primitive task networks of constant generalized partial order width.
Proof.
@@
Let $w$ be the generalized partial order width of the input network, $U$ the set of isolated elements in $T$, and recall that the set $V$ of all remaining elements in $T$ induces a partial order of width at most $w$. First, check whether $p$ is an executable sequence with respect to the preconditions of each action. If not, then return that it is a no-instance. Otherwise, decompose $D_{\prec}[V]$ into $w$ chains $c_{1},\ldots,c_{w}$. This can be done in $\mathcal{O}\!\left(|V|^{2.5}\right)$ time Fellows and McCartin (2003). For $0\leq i\leq|T|$, let $R_{i}[h_{1},\ldots,h_{w}]$ be a boolean variable that is TRUE if and only if the sequence of the first $i$ actions in $p$ is executable in $tn$ such that exactly the first $h_{j}$ tasks are used in each chain $c_{j},j\in[w]$. We compute this variable for all possible assignments in a dynamic programming manner. Initialize all values to FALSE except for $R_{0}[0,\ldots,0]$, which is initialized to TRUE. This completes the calculations for $i=0$. Assume that the calculations for some arbitrary, but fixed $i$ are complete.
@@
Then, update the values for $i+1$ as follows. Iterate over all assignments $R_{i}[h_{1},\ldots,h_{w}]$ that are TRUE. Let $T^{\prime}\subseteq U\cup V$ be such that $t\in T^{\prime}$ if and only if $\alpha(t)=p[i+1]$ and, for each chain $c_{j}$, all the predecessors of $t$ in $\prec^{+}$ in $c_{j}$ are among the first $h_{j}$ tasks in $c_{j}$. Intuitively, $T^{\prime}$ contains the tasks that are executable as the $(i+1)^{\text{th}}$ task in the plan given that the first $h_{j}$ tasks are used in each chain $c_{j}$. Further, let $V^{\prime}\subseteq V$ be such that $t\in V^{\prime}$ if and only if $\alpha(t)=p[i+1]$ and $t$ is among the first $h_{j}$ tasks in its chain $c_{j}$, that is, $V^{\prime}$ contains the already executed tasks of the current action from the chains. Let $o$ denote the number of occurrences of the action $p[i+1]$ in the first $i$ elements of the plan $p$. If $|U\cap T^{\prime}|>o-|V^{\prime}|$, then set $R_{i+1}[h_{1},\ldots,h_{w}]$ to be TRUE. Intuitively, this means that the next task can be taken from $U$ since there is an available task in $U$ that is not already required earlier in the plan. Also, the $(i+1)^{\text{th}}$ task may come from a chain. Thus, for each chain $c_{j},j\in[w],$ such that the $(h_{j}+1)^{\text{th}}$ task is in $T^{\prime}$, set $R_{i+1}[h_{1},\ldots,h_{j-1},h_{j}+1,h_{j+1},\ldots,h_{w}]$ to TRUE. Continue until the computations for $i=|T|$ are completed. Then, the given instance is a yes-instance if and only if there is $R_{|T|}[h_{1},\ldots,h_{w}]$ that is TRUE.
@@
We show by induction over $i$ that the algorithm sets all variables correctly. For $i=0$, the algorithm sets all $R_{0}[h_{1},\ldots,h_{w}]$ correctly as the only way to build an empty plan is to take no tasks from any chain. Now, assume that the algorithm correctly assigned all variables up to some arbitrary, but fixed $i$.
@@
If the algorithm sets some value $R_{i+1}[h_{1},\ldots,h_{w}]$ to TRUE, then we distinguish between two cases. First, assume it does so starting from $R_{i}[h_{1},\ldots,h_{w}]$ being TRUE. Due to the induction hypothesis, there is an executable task sequence that uses the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and whose actions correspond to the first $i$ actions of $p$. Action $p[i+1]$ occurs $o$ times among these. As $|V^{\prime}|$ contains all used tasks with this action from the chains, exactly $o-|V^{\prime}|$ executed tasks of this action are in $U$. As $T^{\prime}\cap U$ contains all executable tasks in $U$, if $|T^{\prime}\cap U|>o-|V|$, then there is a task $t\in U$ that is executable for action $p[i+1]$ and has not already been used. Thus, it is correct to set $R_{i+1}[h_{1},\ldots,h_{w}]$ to TRUE. In the other case, there is $j\in[w]$ such that $R_{i}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},\ldots,h_{w}]$ is TRUE and the $h_{j}^{\text{th}}$ action of the chain $c_{j}$ is executable. Thus, in this case it is also correct to set $R_{i+1}[h_{1},\ldots,h_{w}]$ to TRUE.
@@
For the other direction, suppose there is an executable sequence $r$ of $i$ different tasks in $T$ of the required actions that uses the first $h_{j}$ tasks of each chain $c_{j}$, that is, $R_{i}[h_{1},\ldots,h_{w}]$ should be TRUE. Let the last task in this sequence be $t$ and let $r^{\prime}$ denote the sequence $r$ without $t$. If $t\in U$, then the existence of $r^{\prime}$ implies that $R_{i-1}[h_{1},\ldots,h_{w}]$ is TRUE, and in the computations starting from there we have $|T^{\prime}\cap U|>o-|V|$. If $t\in V$, then $t$ is the $h_{j}^{\text{th}}$ task of a chain $c_{j}$, $R_{i-1}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},\ldots,h_{w}]$ is TRUE, and in the computations starting from there we have $t\in T^{\prime}$. In both cases, the algorithm correctly sets $R_{i}[h_{1},\ldots,h_{w}]$ to TRUE.
@@
We note that the above algorithm is able to also solve the problem on more general networks, specifically even if the isolated vertices in $U$ are allowed to have in-edges from $V$.
Next, we set our sights on State Reachability. While this problem could be solved via a stand-alone argument on antichains in Theorem 8, in the more general setting considered here, we will first need to apply a reduction rule that can be seen as a simpler version of Reduction Rule 1 used for Action Executability.
Reduction Rule R2:
Assume we are given a primitive State Reachability instance. Let $a_{1}$ and $a_{2}$ be two equivalent actions. For all $t\in T$ such that $\alpha(t)=a_{2}$, set $\alpha(t)=a_{1}$.
Observation 12.
Reduction Rule R2 is safe.
Proof.
Assume there is a task sequence $t_{1},\ldots,t_{\ell}$ solving the original instance. Then, the same sequence solves the reduced instance as it gives the same path in the state transition graph in either instance. The same holds for the other direction. ∎
With this, we are now ready to also establish tractability for State Reachability.
Theorem 13.
State Reachability is polynomial-time solvable on primitive task networks with constantly many states and constant generalized partial order width.
Proof.
Let $k$ be the number of states and $w$ the generalized partial order width of the input instance. Set $U$ to be the set of isolated elements in $T$, and recall that the set $V$ of all remaining elements in $T$ induces a partial order of width at most $w$. We begin by exhaustively applying Reduction Rule R2, which can be done in time at most ${\mathcal{O}\!\left(|{\cal I}|^{3}k^{2}\right)}$. Afterwards, we note that there is at most one action for each of the at most $(k+1)^{k}$ equivalence classes.
The algorithm now decomposes $D_{\prec}[V]$ into $w$ chains $c_{1},\ldots,c_{w}$. This can be done in $\mathcal{O}\!\left(|V|^{2.5}\right)$ time Fellows and McCartin (2003). Let $A={\{a_{1},\ldots,a_{|A|}\}}$. For each state $s$ in the state transition graph, let $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{|A|}]$ be a boolean variable that is TRUE if and only if there is an executable task sequence in $tn$ starting from $s_{0}$ and ending in state $s$ that employs exactly the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and contains exactly $u_{i}$ tasks with action $a_{i}$, for all $i\in[|A|]$.
We compute this variable for all possible assignments in a dynamic programming manner. Initialize all values to FALSE except for $R_{s_{0}}[0,\ldots,0][0,\ldots,0]$, which is initialized to TRUE. Whenever a variable $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{|A|}]$ that was FALSE before is set to TRUE, also perform the following update step. Let $T^{\prime}\subseteq T$ be such that $t\in T^{\prime}$ if, for each chain $c_{j}$, all the predecessors of $t$ in $\prec^{+}$ in $c_{j}$ are among the first $h_{j}$ tasks in $c_{j}$. Consider all the states $s^{\prime}$ that are adjacent to $s$ in ${\cal G}$. If there is $j\in[w]$ such that the $(h_{j}+1)^{\text{th}}$ task of the chain $c_{j}$ is in $T^{\prime}$, has action $a_{i}$, and $a_{i}$ connects $s$ and $s^{\prime}$ in ${\cal G}$, then set $R_{s^{\prime}}[h_{1},\ldots,h_{j-1},h_{j}+1,h_{j+1},\ldots,h_{w}][u_{1},\ldots,u_{i-1},u_{i}+1,u_{i+1},\ldots,u_{|A|}]$ to TRUE. This corresponds to taking the next task from a chain and adding it to a given sequence. Another update rule corresponds to adding a task from $U$ instead. If there is an action $a_{i}$ that connects $s$ and $s^{\prime}$ in ${\cal G}$, then define the following. Let $T_{i}={\{t\in T\mid\alpha(t)=a_{i}\}}$ and let $v_{i}$ denote the number of tasks in $V\cap T_{i}$ that are among the first $h_{j}$ tasks of a chain $c_{j}$. If $u_{i}-v_{i}<|T_{i}\cap T^{\prime}\cap U|$, then set $R_{s^{\prime}}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{i-1},u_{i}+1,u_{i+1},\ldots,u_{|A|}]$ to TRUE. Once no more variables are updated and set to TRUE, the computation is complete. Then, a given State Reachability instance is a yes-instance if and only if there is $R_{s_{g}}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{w}]$ that is TRUE. A given Action Executability instance is a yes-instance if and only if there is $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{w}]$ that is TRUE such that, for each action $a_{i}\in S$, we have $\min(a)\geq u_{i}$.
We show that the algorithm sets all variables correctly by an inductive argument over $\sum_{i=1}^{|A|}u_{i}$. Note that when one TRUE variable sets another variable to TRUE, this increases this sum by exactly 1. For $\sum_{i=1}^{|A|}u_{i}=0$, the algorithm sets all variables correctly as the only way to build an empty plan is to take no tasks from any chain and remain in the initial state, and thus, the only TRUE variable in this case is $R_{s_{0}}[0,\ldots,0][0,\ldots,0]$. Now, assume that the algorithm correctly assigned all variables up to some arbitrary, but fixed $\sum_{i=1}^{|A|}u_{i}$.
If some value $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{w}]$ is set to TRUE by the algorithm, then we distinguish between two cases. First, assume it does so starting from some $R_{s^{\prime}}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{i-1},u_{i}-1,u_{i+1},\ldots,u_{|A|}]$ being TRUE with $i\in[|A|]$. Then, $a_{i}$ is executable in state $s^{\prime}$ and transforms it into $s$. Further, due to the induction hypothesis, there is an executable task sequence that uses the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and consists of $u_{\ell}$ tasks of action $a_{\ell}$, for each $\ell\in[|A|]\setminus{\{i\}}$, and $u_{i}-1$ tasks of action $a_{i}$. Thus, it contains $u_{i}-v_{i}$ tasks in $U$ with action $a_{i}$. There are more than $u_{i}-v_{i}$ executable tasks of action $a_{i}$ in $U$, and thus, one of the remaining tasks of action $a_{i}$ can be appended to the sequence. The resulting sequence is executable in $tn$ starting from $s_{0}$, uses the same number of tasks of each chain and action, except one more task with action $a_{i}$. Thus, it is correct to set $R_{s^{\prime}}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{w}]$ to TRUE. In the other case, it sets the variable to TRUE starting from some $R_{s^{\prime}}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},h_{w}][u_{1},\ldots,u_{i-1},u_{i}-1,u_{i+1},\ldots,u_{|A|}]$ being TRUE with $j\in[w]$, $i\in[|A|]$. Then, due to the induction hypothesis, there is an executable task sequence that uses the first $h_{\ell^{\prime}}$ tasks of each chain $c_{\ell^{\prime}}$, for each $\ell^{\prime}\in[w]\setminus\{j\}$, but only uses the first $h_{j}-1$ tasks of the chain $c_{j}$, and it consists of $u_{\ell}$ tasks of action $a_{\ell}$, for each $\ell\in[|A|]\setminus{\{i\}}$, and $u_{i}-1$ tasks of action $a_{i}$. Further, the $h_{j}^{\text{th}}$ task of chain $h_{j}$ has action $a_{i}$, is executable at the end of this sequence, and transforms state $s^{\prime}$ into $s$. Thus, in this case it is also correct to set the variable to TRUE.
For the other direction, suppose there is an executable sequence $r$ of $\sum_{i=1}^{|A|}u_{i}$ different tasks in $T$ such that, for each $\ell\in[|A|]$, there are $u_{\ell}$ tasks with action $a_{\ell}$ that use the first $h_{j}$ tasks of each chain $c_{j}$. Suppose $r$ ends in state $s$, that is, $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{|A|}]$ should be TRUE. Let the last task in this sequence be $t$ with $\alpha(t)=a_{i}$ and let $r^{\prime}$ denote the sequence $r$ without $t$. Let $r^{\prime}$ end in state $s^{\prime}$. If $t\in U$, then the existence of $r^{\prime}$ implies $R_{s^{\prime}}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{i-1},u_{i}-1,u_{i+1},\ldots,u_{|A|}]$ is TRUE. In the update step of this variable, $t$ is executable, and thus, $u_{i}-v_{i}<|T_{i}\cap T^{\prime}\cap U|$. Then, the algorithm correctly sets $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{|A|}]$ to TRUE. If $t\in V$, then $t$ is the $h_{j}^{\text{th}}$ task of a chain $c_{j}$, $R_{s^{\prime}}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},\ldots,h_{w}][u_{1},\ldots,u_{i-1},u_{i}-1,u_{i+1},\ldots,u_{|A|}]$ is TRUE, and in the computations starting from there $t\in T^{\prime}$. In both cases, the algorithm correctly sets the variable to TRUE.
For each variable that is TRUE, we have that $s$ is among the $k$ states in ${\cal G}$, $0\leq h_{j}\leq|T|$ for $j\in[w]$, and $0\leq\sum_{i=1}^{|A|}u_{i}\leq|T|$. Thus, at most ${\mathcal{O}\!\left(|T|^{w}\cdot k\cdot|T|^{|A|}\right)}$ variables are considered. For each of these, computing $T^{\prime}$, the $u_{i}$, and $v_{i}$ takes $\mathcal{O}\!\left(|T|+|\prec|\right)$ time, it can update at most $(w+1)\cdot|A|$ variables, and it can be checked in $\mathcal{O}\!\left(|A|\right)$ time if it yields a yes-instance. Hence, the algorithm takes ${\mathcal{O}\!\left(|T|^{w+|A|}\cdot k\cdot(|T|+|\prec|+(w+1)\cdot|A|)\right)}={\mathcal{O}\!\left(|T|^{w+|A|}\cdot k\cdot(|T|+|\prec|+w|A|)\right)}$ time. Once again, the theorem now follows by the bound on $|A|$ established in the first paragraph of the proof. ∎
We observe that, once more, the above algorithm is actually able to solve the more general problem where the isolated vertices in $U$ are allowed to have in-edges from $V$. Finally, we note that Reduction Rule R1 by itself does not suffice to transfer this result to Action Executability as the number of neighborhoods of tasks is not bounded (as it was for Theorem 10). Nevertheless, the desired result can be obtained by carefully adapting the dynamic program from Theorem 13 and incorporating the idea of the reduction rule on the (isolated) tasks in $U$.
Theorem 14.
Action Executability is polynomial-time solvable on primitive task networks with constantly many states and constant generalized partial order width.
Proof.

### v2

@@ Appendix A Appendix: Proof Details
Let $\max_{r}(a)$ and $\min_{r}(a)$ be defined as above, but for the reduced instance. Hence, $\max_{r}(a_{1})=\max(a_{1})+\max(a_{2})$ and $\min_{r}(a_{1})=\min(a_{1})+\min(a_{2})$. Thus, any task sequence that solves the original instance also solves the reduced instance. Now assume there is a task sequence that solves the reduced instance. If the same task sequence solves the original instance, we are done. Otherwise, let $x_{1}$ and $x_{2}$ be the number of occurrences of $a_{1}$- and $a_{2}$-tasks, respectively, in the sequence resolving the reduced instance according to the original mapping $\alpha$. Note that $\max_{r}(a_{1})=\max(a_{1})+\max(a_{2})\geq x_{1}+x_{2}\geq\min(a_{1})+\min(a_{2})$. Due to the equivalence in their actions and neighborhoods in $\prec^{+}$, $a_{1}$- and $a_{2}$-tasks can be used interchangeably in the sequence. Thus, with the above inequalities, we can find a valid solution for the original instance, i.e., one in which $\max(a_{1})\geq x_{1}\geq\min(a_{1})$ and $\max(a_{2})\geq x_{2}\geq\min(a_{2})$. ∎
@@
Note that an instance of Action Executability with $|S|>|T|$ is a trivial no-instance, so we assume $|S|\leq|T|$.
Proof of Theorem 4.
@@
We begin by exhaustively applying Reduction Rules R0 and R1, which takes at most $\mathcal{O}(|{\cal I}|^{3})$ time. Afterwards, we have either correctly solved the instance, or—since the network is an antichain—obtained a new equivalent instance where the number of actions is bounded by a function of the number $k$ of states, in particular $|A|\leq(k+1)^{k}$.
@@
The algorithm now branches over the at most $2^{k}\cdot|A|^{k}$ simple edge paths in ${\cal G}$ that start at $s_{0}$. Let $C$ be the set of distinct simple edge cycles in ${\cal G}$, where two cycles are distinct if and only if they do not contain the same vertices. Then, $|C|\leq 2^{k}\cdot|A|^{k}$. The algorithm branches on all possible subsets $C^{\prime}\subseteq C$ to decide which of these cycles are used at least once by a solution sequence. To ensure there is a walk, set all cycles in $C^{\prime}$ as unmarked and exhaustively mark cycles that share a vertex with the chosen path or a marked cycle. Proceed with $C^{\prime}$ only if all cycles are marked by this procedure.
@@
We write an Integer Linear Program (ILP) as follows. Let the ILP have a variable $x_{c}$ for every $c\in C^{\prime}$. For each action $a$ and $c\in C^{\prime}$, let $c(a)$ denote the number of occurrences of $a$ in $c$. Further, let $p(a)$ denote the number of occurrences of $a$ in the chosen path. For each $a\in A$, add constraints to require
$\min(a)\leq p(a)+c_{1}(a)x_{c_{1}}+\dots c_{|C^{\prime}|}(a)x_{c_{|C^{\prime}|}}$ $\leq\max(a),$
@@
where $c_{1},\ldots,c_{|C^{\prime}|}$ are the elements of $C^{\prime}$. For all $c\in C^{\prime}$, add a constraint $x_{c}\geq 1$ to ensure each of the selected cycles is walked at least once. The given instance is a yes-instance if and only if there is a solution to the ILP.
@@
Computing $\min(a)$ and $\max(a)$ for all $a\in A$ takes $\mathcal{O}\!\left(|T|+|S|\right)$ time. Combining the choices for the path and $C^{\prime}$ gives at most $2^{k}\cdot|A|^{k}\cdot 2^{2^{k}\cdot|A|^{k}}$ branches. Testing a set $C^{\prime}$ by the marking procedure takes $\mathcal{O}\!\left(|C|^{2}k\right)$ time, which is dominated by the time to solve the ILP. The ILP has at most $|C|\leq 2^{k}\cdot|A|^{k}$ variables and $2|A|+|C|={\mathcal{O}\!\left(2^{k}\cdot|A|^{k}\right)}$ constraints, in which the absolute values of all coefficients are bounded by $|T|$. By the result of Lenstra [1983] and its subsequent improvements Frank and Tardos [1987]; Kannan [1987] an ILP instance ${\cal I}_{\text{ILP}}$ with $n$ variables can be solved in $\mathcal{O}\!\left(n^{O(n)}\cdot|{\cal I}_{\text{ILP}}|\right)$ time. We note that, for the given ILP, $|I_{\text{ILP}}|={\mathcal{O}\!\left(2^{k}\cdot|A|^{k}\cdot\log(|T|)\right)}$. Thus, solving the ILP takes $\mathcal{O}\!\left((2^{k}\cdot|A|^{k})^{O(2^{k}\cdot|A|^{k})}\cdot 2^{2k}\cdot|A|^{2k}\cdot\log(|T|)\right)$ time. Hence, the total runtime can be upper-bounded by
@@
Recalling the bound on $|A|$ in the first paragraph of the proof, we obtain that any primitive instance of Action Executability with $\prec^{+}=\emptyset$ can be solved in time $\mathcal{O}(f(k)\cdot\log(|{\cal I}|)+|{\cal I}|^{3})$, for some computable function $f$. ∎
Proof of Theorem 5.
@@
Let $w$ be the generalized partial order width of the input network, $U$ the set of isolated elements in $T$, and recall that the set $V$ of all remaining elements in $T$ induces a partial order of width at most $w$. First, check whether $p$ is executable from $s_{0}$. If not, then return that it is a no-instance. Otherwise, decompose $D_{\prec}[V]$ into $w$ chains $c_{1},\ldots,c_{w}$ in $\mathcal{O}\!\left(|V|^{2.5}\right)$ time Fellows and McCartin [2003]. For $0\leq i\leq|T|$, let $R_{i}[h_{1},\ldots,h_{w}]$ be a boolean variable that is TRUE if and only if the sequence of the first $i$ actions in $p$ is executable in $tn$ such that exactly the first $h_{j}$ tasks are used in each chain $c_{j},j\in[w]$. We compute this variable for all possible assignments via dynamic programming. Initialize all values to FALSE except for $R_{0}[0,\ldots,0]$, which is TRUE. This completes the calculations for $i=0$. Assume that the calculations for some arbitrary, but fixed $i$ are complete.
@@
Then, update the values for $i+1$ as follows. Iterate over all assignments $R_{i}[h_{1},\ldots,h_{w}]$ that are TRUE. Let $T^{\prime}\subseteq U\cup V$ be such that $t\in T^{\prime}$ if and only if $\alpha(t)=p[i+1]$ and, for each chain $c_{j}$, all the predecessors of $t$ in $\prec^{+}$ in $c_{j}$ are among the first $h_{j}$ tasks in $c_{j}$. Intuitively, $T^{\prime}$ contains the tasks that match the $(i+1)^{\text{th}}$ action in the plan and have that each of their predecessors in $\prec^{+}$ is among the first $h_{j}$ tasks of a chain $c_{j}$. Further, let $V^{\prime}\subseteq V$ be such that $t\in V^{\prime}$ if and only if $\alpha(t)=p[i+1]$ and $t$ is among the first $h_{j}$ tasks in its chain $c_{j}$, that is, $V^{\prime}$ contains the already executed tasks of the current action from the chains. Let $o$ denote the number of occurrences of the action $p[i+1]$ in the first $i$ elements of the plan $p$. If $|U\cap T^{\prime}|>o-|V^{\prime}|$, then set $R_{i+1}[h_{1},\ldots,h_{w}]$ to be TRUE. Intuitively, this means that the next task can be taken from $U$ since there is an available task in $U$ that is not already required earlier in the plan. Also, the $(i+1)^{\text{th}}$ task may come from a chain. Thus, for each chain $c_{j},j\in[w],$ such that the $(h_{j}+1)^{\text{th}}$ task is in $T^{\prime}$, set $R_{i+1}[h_{1},\ldots,h_{j-1},h_{j}+1,h_{j+1},\ldots,h_{w}]$ to TRUE. Continue until the computations for $i=|T|$ are completed. Then, the given instance is a yes-instance if and only if there is $R_{|T|}[h_{1},\ldots,h_{w}]$ that is TRUE.
@@
We show by induction over $i$ that the algorithm sets all variables correctly. For $i=0$, the algorithm sets all $R_{0}[h_{1},\ldots,h_{w}]$ correctly as the only way to build an empty plan is to use no tasks at all. Now, assume that the algorithm correctly assigned all variables up to an arbitrary, but fixed $i$.
@@
If the algorithm sets some value $R_{i+1}[h_{1},\ldots,h_{w}]$ to TRUE, then we distinguish between two cases. First, assume it does so starting from $R_{i}[h_{1},\ldots,h_{w}]$ being TRUE. Due to the induction hypothesis, there is an task sequence respecting $\prec^{+}$ that uses the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and whose actions correspond to the first $i$ actions of $p$. Action $p[i+1]$ occurs $o$ times among these. As $|V^{\prime}|$ contains all used tasks with this action from the chains, exactly $o-|V^{\prime}|$ executed tasks of this action are in $U$. As $T^{\prime}\cap U$ contains all executable tasks in $U$, if $|T^{\prime}\cap U|>o-|V|$, then there is a task $t\in U$ that is executable for action $p[i+1]$ and has not already been used. Thus, it is correct to set $R_{i+1}[h_{1},\ldots,h_{w}]$ to TRUE. In the other case, there is $j\in[w]$ such that $R_{i}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},\ldots,h_{w}]$ is TRUE and the $h_{j}^{\text{th}}$ action of the chain $c_{j}$ is executable. Thus, in this case it is also correct to set $R_{i+1}[h_{1},\ldots,h_{w}]$ to TRUE.
@@
For the other direction, suppose there is a sequence $\bar{t}$ of $i$ distinct tasks in $T$ of the required actions that uses the first $h_{j}$ tasks of each chain $c_{j}$, that is, $R_{i}[h_{1},\ldots,h_{w}]$ should be TRUE. Let the last task in this sequence be $t$ and let $\bar{t}^{\prime}$ denote the sequence $\bar{t}$ without $t$. If $t\in U$, then the existence of $\bar{t}^{\prime}$ implies that $R_{i-1}[h_{1},\ldots,h_{w}]$ is TRUE, and in the computations starting from there we have $|T^{\prime}\cap U|>o-|V|$. If $t\in V$, then $t$ is the $h_{j}^{\text{th}}$ task of a chain $c_{j}$, $R_{i-1}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},\ldots,h_{w}]$ is TRUE, and in the computations starting from there we have $t\in T^{\prime}$. In both cases, the algorithm correctly sets $R_{i}[h_{1},\ldots,h_{w}]$ to TRUE.
@@
Proof of Theorem 6.